OpenStudy (baldymcgee6):

Verify: x^2 + y^2 = 1 Take the derivative of both sides 2x + 2y*y' = 0 y' = -x/y Find the second derivative y'' = (-y + xy')/y^2

OpenStudy (anonymous):

I think u r correct...

OpenStudy (baldymcgee6):

Wolfram says y'' = -1/y

OpenStudy (anonymous):

y'' = (-y + xy')/y^2 Wolfram says y'' = -1/y Substitute y' = -x/y y'' = (-y + x(-x/y))/y^2 y'' = -(x^2+y^2)/y^3 substitute x^2 = 1-y^2 simplify, y'' = -1/y^3 is what I got . . . :-/

OpenStudy (baldymcgee6):

hmmm

OpenStudy (anonymous):

I'll try it again from scratch and see what happens

OpenStudy (baldymcgee6):

I tried that as well... Wolfram's explanation: y' = -x/y derivative of x =1 y'' = -1/y

OpenStudy (baldymcgee6):

what if we went: y' = -xy^-1

OpenStudy (anonymous):

Should be the same if you use product rule.

OpenStudy (baldymcgee6):

Yeah, i'm sure that the original y'' I found was correct, but it just bugs me how they simplified it so nicely.

OpenStudy (anonymous):

Pff, yeah, beats me.

OpenStudy (baldymcgee6):

Yeah, thanks for the help, i'll leave it up to see if anyone else has any bright ideas.

OpenStudy (baldymcgee6):

@satellite73, @Hero, @AccessDenied, @hartnn, @lgbasallote

hartnn (hartnn):

wolfram is showing me correctly as -x/y

OpenStudy (anonymous):

What about second derivative, @hartnn ?

OpenStudy (baldymcgee6):

@hartnn the first derivative of x^2 + y^2 =1 is -x/y, but wolfram somehow simplified the second derivative into -1/y